how much water must be evaporated from 32 gallons of a 10% salt solution in order to obtain a 20% salt solution
Answer by richwmiller(17219)

0.1*32=0.2*x
3.2=0.2x
x=16 new total amount
Remove 16
0.2*16=3.2
You want to double the concentration so you half the amount of water